A rectangular lawn has sides which measure x − 5 feet and 3x feet. The perimeter is 62 feet. Find the length of each side of the lawn.
Answer by lwsshak3(11628)

A rectangular lawn has sides which measure x − 5 feet and 3x feet. The perimeter is 62 feet. Find the length of each side of the lawn.
***
let 3x=length of rectangular lawn
let (x-5)=width of rectangular lawn
2*length+2*width=perimeter
2*3x+2(x-5)=62
6x+2x-10=62
8x=72
x=9
3x=27
x-5=4
length of rectangular lawn=27 ft
width of rectangular lawn=4 ft